OpenStudy (anonymous):

The point (5, -1) is a solution of the equation: y = 2x − 11 True False

OpenStudy (anonymous):

Just plug it in and see. y = 2x - 11 5 is x and y is - 1. -1 = 2(5) - 11 -1 = 10 - 11 -1 = -1 So, yes, it is true.

OpenStudy (anonymous):

thank you (: